# Algorithm for finding local integrals of motion in quantum lattice models in the thermodynamic limit

### v1

@@ Abstract
Concluding remarks – We have derived and implemented a numerical approach that provides exact translationally-invariant LIOMs and the corresponding Mazur bounds for infinite quantum systems for which the dynamics is set by either local Hamiltonian or by local unitary transformations. We are able to obtain the exact LIOMs because our approach, similarly to Refs. [74, 69], is targeting the short-time dynamics (as encoded, e.g., in the commutators $[\hat{O},\hat{H}]$) under which local operator $\hat{O}$ remains local. It is an important advantage over approaches focused on the infinite-time properties that are encoded in the diagonal matrix elements of observables calculated in the basis of the eigenstates of the Hamiltonian. In the latter case, the finite-size effects seem unavoidable.
Like all numerical methods, our approach also has limitations. This bottleneck concerns the number of local observables, $N_{O}\sim 10^{4}-10^{5}$, from which LIOMs are built by solving the eigenproblem in Eq. (1). In case of spin systems, we still considered all operators supported on up to 10 consecutive sites. On the one hand, one may expect that finding very complex LIOMs may be problematic. On the other hand, such LIOMs only have a limited influence on the dynamics of physically relevant observables, which are typically represented by few-body local operators.
@@
Our approach is not based on diagonalization of the Hamiltonian or a unitary matrix. Therefore, it can be applied to problems that are hardly accessible to other numerical methods, e.g., when one studies problems beyond one-dimensional chains. In particular, it also identifies quasilocal LIOMs as well as LIOMs in the Hubbard model. Furthermore, we have demonstrated that this approach also applies to nearly integrable systems, where it singles out slowly decaying observables and provides reasonable estimates of the corresponding relaxation rates. It is rather obvious that one may also carry out the same calculations for systems that are not translationally invariant. In the latter case, although one cannot study infinite systems anymore, the accessible system sizes are still much larger than is the case within approaches that are based on the diagonalization procedure.
@@
Acknowledgments. J.P. acknowledges support by the National Science Centre (NCN), Poland via project 2023/49/N/ST3/01033. M.M. acknowledges support by the National Science Centre (NCN), Poland via project 2020/37/B/ST3/00020.

### v2

@@ IV Concluding remarks
We have derived and implemented a general numerical approach that provides exact translationally-invariant LIOMs and the corresponding Mazur bounds for infinite quantum systems for which the dynamics is set by either local Hamiltonian or by local unitary transformations. We can obtain the exact LIOMs for the thermodynamic limit because our approach, similarly to Refs. [69, 70, 71], is targeting the short-time dynamics (as encoded, e.g., in the commutators $[\hat{O},\hat{H}]$) under which local operator $\hat{O}$ remains local. Consequently, our results have an important advantage over approaches focused on the infinite-time properties, which are encoded in the diagonal matrix elements of observables calculated in the basis of the Hamiltonian’s eigenstates. In the latter case, the thermodynamic limit is unreachable, and finite-size effects seem unavoidable.
@@
Diagonalization of the Hamiltonian (exact or approximate) is a common bottleneck for most problems studied in the context of interacting quantum systems. However, our approach is not based on the diagonalization of the Hamiltonian or a unitary matrix. Therefore, it can be applied to problems that are hardly accessible to other numerical methods. In particular, it enables the study of problems beyond one-dimensional chains. It also identifies quasilocal conserved quantities in the XXZ chain and LIOMs in much more demanding system, i.e., in the Hubbard model. Furthermore, we have demonstrated that this approach also applies to nearly integrable systems, where it singles out slowly decaying observables and provides reasonable estimates of the corresponding relaxation rates. Studying the perturbed XXZ chain, we have found that these slow modes are not inherited from a single integrable Hamiltonian but rather from LIOMs which exist in a broader class of integrable models.
It is rather obvious that one may also carry out similar calculations for systems that are not translationally invariant, e.g., containing inhomogeneous potentials. In the latter case, although one cannot study infinite systems anymore, the accessible system sizes are still much larger than those in approaches based on the diagonalization procedure.
Like all numerical methods, our approach also has limitations. This bottleneck concerns the number of local observables, $N_{O}\sim 10^{4}-10^{5}$, from which LIOMs are built by solving the eigenproblem in (1). In the case of spin systems, we still considered all operators supported on up to 10 consecutive sites. On the one hand, one may expect that finding very complex LIOMs may be problematic. On the other hand, such LIOMs only have a limited influence on the dynamics of physically relevant observables, which are typically represented by few-body local operators.
Acknowledgements.
@@
We acknowledge fruitful discussions with Marko Žnidarič and Vir B. Bulchandani. J.P. acknowledges support by the National Science Centre (NCN), Poland via project 2023/49/N/ST3/01033. M.M. acknowledges support by the National Science Centre (NCN), Poland via project 2020/37/B/ST3/00020.